OpenStudy (anonymous):

What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3? A.243 B.405 C.1,215

OpenStudy (asnaseer):

you just asked and closed this question here: http://openstudy.com/updates/4fe6279ee4b02c91101afdf1

OpenStudy (asnaseer):

please don't close and repost the same questions - it is considered spamming and may lead to suspension.

OpenStudy (zepp):

This is a geometric sequence with first term 5; common ratio 3. Mean that 1st term 2nd term = 1st term * 3 3rd term = 2nd term * 3 ... and so on.

OpenStudy (zepp):

And to find any term, use the rule \(x = ar^{n-1}\) where n is the term you are trying to find.

OpenStudy (campbell_st):

a term in a geometric series is found using \[T _{n} = ar^{n -1}\] you have a = 5, r = 3 and n = 5 \[T _{5} = 5\times 3^{5 -1}\] just evaluate

OpenStudy (zepp):

We are looking for the 5th term, plug replace a by 5; r by 3; n by 5 and solve for x; \(\large x = 5*3^{5-1}=5*3^4=405\)